A chemical company has two varietys of toxic chemical, a 10% solution and a 20% solution. A customer wants 10 liters of a 16% solution. How many liters of each concentration should be mixed to fill the order?
Tried all sorts of different ways and just can't seem to get this one right. Not by using the example my teacher gave me. Any help would be greatly appreciated.

Answer by checkley77(12844)

.20x+.10(10-x)=.16*10
.20x+1-.10x=1.6
.10x=1.6-1
.10x=.6
x=.6//.10
x=6 liters of 20% is needed.
Proof:
.20*6+.10(10-6)=1.6
1.2+.10*4=1.6
1.2+.4=1.6
1.6=1.6